# Observation of the Tetragonal-Orthorhombic transition in polycrystalline YBa2Cu3O6+δ samples during the oxidation process at constant temperature via thermogravimetry and differential thermometry analysis

### v1

@@ IV Conclusions
However, considering that the Tetragonal phase advances within the Orthorhombic zone of the material only in the presence of a short-range order superstructure (O-II), metastability should not occur for temperatures above 400 oC, where the Orthorhombic phase is purely long-range order (O-I). In that sense, the appearance of metastability in this transition for the data measured and shown here at TO = 691 oC is puzzling. We rule out that this is due to the appearance of other superstructures already measured in the material (O-III, O-V, O-VIII or even O-II) as these occur at lower temperature values[8]. Despite this, it is also noteworthy that for TO values above 600 oC, the T-O transition occurs for $\delta$ values in coherence with the material’s Oxygen saturation value. Additionally, saturation rates for TO values above 600 oC are not only the highest but also similar to each other (difference in the m value for the tOxyg shown in Fig. 2). We observed that for these TO values, Oxygen saturation occurs at very short times, and the final values of $\delta$ reached maintain a proportional relationship with the $\delta$ values at which the T-O transition occurs. This indicates that the metastability at these TO values depends not only on defects but also on the abrupt rates (for structural changes) at which oxidation occurs. Considering this possibility, we propose that metastability of the Tetragonal phase at these higher TO values (incursing into the O-I superstructure) and under our experimental conditions, occurs due to the violent oxygenation rate. We understand it could be useful for confirming this assertion, to extend these studies by conducting oxidation experiments at lower rates, as we notice that this could bring the measured results of the T-O transition closer to the theoretically predicted line (Fig. 6). For experimental techniques we have access in our laboratories, it was not possible to perform such experiments since a partially saturated O2 atmosphere implies working with a mixture of both gases (N2 and O2), driving the experiment unstable and noisy.
V Final remarks
@@
From YBCO powder obtained by mechanical milling of superconducting pellets, both mass evolution and the value of the relative temperature referenced to an inert sample were studied as functions of the oxidation time. Starting from completely deoxygenated samples, a maximum oxidation of $\delta$ = 0.97 was achieved for TO = 394 oC. It was also observed that, for short times, tOxyg = 5 min, the material reaches maximum oxidation for TO = 691 oC, saturating within 5 % of tolerance in the m value. Additionally, the values of differential thermometry indicated the oxidation process as an exothermic type. It was also observed that, associated with this process, there is a second exothermic process, which we link to the T-O transition in the material. Both processes were well-differentiated from each other, appearing one order of magnitude apart in a tOxyg scale measured in units of seconds. Subsequently, a $\delta$ value was associated with the T-O transition for each TO. With these values, a structural phase diagram for the material was constructed, observing not only the T-O phase transition but also the interface between O-I and O-II superstructures in the Orthorhombic phase. Finally, based on variations of our results compared to theoretical models predicting the structural phase diagram of the material[6, 8], different mechanisms were proposed that generate metastability during the T-O transition in the oxidation process. On one hand, at low temperatures, a mechanism of multiple nucleation centers of a superstructure with short-range order (O-II) was proposed to maintain the metastability of the Tetragonal phase. On the other hand, at high temperatures, a mechanism where the high speed in the oxidation process plays a very important role was proposed. Both cases are accompanied by a significant density of structural defects intrinsic to the nature of the samples with which we work. In addition to this, the enormous potential of a thermogravimetry setup combined with in-situ differential thermal analysis was highlighted. With these simple and accessible results, it was possible to replicate results previously obtained in large facilities using techniques such as neutron diffraction and high-energy X-ray diffraction[14].
@@ VI Acknowledgments
This work was partially supported by MINCyT throw the PICT-2017-2898 FONCyT project. LG performed the TGA-DTA experiments and oxygenation analysis. CES performed the YBCO pellet growth and X-ray experiments and analysis. JAM contributed to the YBCO pellet growth, TGA-DTA experiments and X-ray experiments. RFL performed the TGA-DTA experiments and structural transition analysis, conducted the research and wrote out the manuscript. All authors contributed equally to discussion of ideas and correction of the manuscript.

### v2

@@ IV Discussion
On the other hand, at high TO values conditions are others. For temperatures above 400 oC, the Orthorhombic phase is purely long-range order (O-I), with no presence of other superstructures (O-III, O-V, O-VIII or even O-II) since they exist at much lower temperature values[8]. However, it is also noteworthy that above 600 oC, relative $\delta$ values associated to the T-O transition (Fig. 4) are in coherence with relative $\delta$ values associated to the Oxygen saturation of the material (Fig. 2). Additionally, rates of Oxygen saturation above 600 oC are not only the highest but also similar between them (Fig. 2), with very short times of saturation. This indicates that metastability at high TO values depends not only on defects but also on the abrupt rates at which oxidation occurs (abrupt even for structural changes). Considering this possibility, we propose that discrepancies between our data and the theoretically predicted ones for the TO transition values above 600 oC, occurs due to the violent oxygenation rate that the material is exposed, incursing into the O-I superstructure in a metastable condition.
V Conclusions
@@
From YBCO powder obtained by mechanical milling of superconducting pellets, both mass evolution and differential temperature referenced to an inert sample were studied as functions of the oxidation time. Starting from completely deoxygenated samples, a maximum oxidation of $\delta$ = 0.97 was achieved for TO = 394 oC. It was also observed that, for short times, tOxyg = 5 min, the material reaches maximum oxidation for TO = 691 oC, saturating within 5 % of tolerance in the m value. Additionally, the values of differential temperature indicated the oxidation process as an exothermic type. It was also observed that, associated with this process, there is a second exothermic process very well differentiated, which we link to the T-O transition in the material. Subsequently, a $\delta$ value was associated with the T-O transition for each TO. With these values, a structural phase diagram for the material was constructed, observing not only the T-O phase transition but also the interface between O-I and O-II superstructures in the Orthorhombic phase. Finally, based on variations of our results compared to theoretical models predicting the structural phase diagram of the material[6, 8], different mechanisms were proposed that generate metastability during the T-O transition in the oxidation process. On one hand, at low temperatures, a mechanism of multiple nucleation centers of a superstructure with short-range order (O-II) was proposed to maintain the Tetragonal phase metastably into the Orthorhombic one. On the other hand, at high temperatures, a mechanism where the high speed in the oxidation process plays a very important role was proposed. Both cases are accompanied by a significant density of structural defects intrinsic to the nature of the samples with which we work. In addition to this, the enormous potential of a thermogravimetry setup combined with in-situ differential thermal analysis was highlighted. With these simple and accessible results, it was possible to replicate results previously obtained in large facilities using techniques such as neutron diffraction and high-energy X-ray diffraction[14].
@@ VI Acknowledgments
This work was partially supported by MINCyT through the PICT-2017-2898 FONCyT project. LG performed the TGA-DTA experiments and oxygenation analysis. CES performed the YBCO pellet growth and X-ray experiments and analysis. JAM contributed to the YBCO pellet growth, TGA-DTA experiments and X-ray experiments. RFL performed the TGA-DTA experiments and structural transition analysis, conducted the research and wrote out the manuscript. All authors contributed equally to discussion of ideas and correction of the manuscript.